How much charge is represented by 4,600 electrons?
1
Expert's answer
2020-09-28T08:11:29-0400

We can find the charge from the formula:


"Q = Ne,"

here, "N = 4600" is the quantity of the electrons, "e = 1.6 \cdot 10^{-19} \ C" is the charge of the electron.

Then, we get:


"Q = 4600 \cdot 1.6 \cdot 10^{-19} \ C = 7.36 \cdot 10^{-19} \ C."

Answer:

"Q = 7.36 \cdot 10^{-19} \ C."